I need my answer checked. Thank you! My answers in bold.

To recruit homeless participants for a study you post fliers at all the homeless shelters in the city describing your study and asking people to contact you if they are interested in participating. You should be concerned about

survivorship bias

self-selection bias

construct validity

face validity

An IQ test has a mean of 100 and a standard deviation of 15 in the population. What is the z score for someone with an IQ of 70

2.2

-2

-1.5

0

IQ is normally distributed with a population mean of 100 and standard deviation of 15. What proportion of people have an IQ above 110?

.2525

.7475

.1587

.2475
Answer #1

Solution :

a)

Self-selection bias

(b)

mean = \mu = 100

standard deviation = \sigma = 15

x = 70

z = (x - \mu ) / \sigma = (70 - 100) / 30 = -30 / 15 = -2

z score = -2

c)

mean = \mu = 100

standard deviation = \sigma = 15

P(x > 110) = 1 - P(x < 110)

= 1 - P((x - \mu ) / \sigma < (110 - 100) / 15)

= 1 - P(z < 0.6667)

= 1 - 0.7475   

= 0.2525

Probability = 0.2525